Assume a competitive firm faces a market price of $100, a cost curve of C 0.64q 25q+1,600 and a marginal cost curve of MC 1.28q 25. The firms profit maximizing output level is 58.59 units, the profit per unit is $10.19, and total profit is: $597.03 However, if the firm wanted to maximize the profit per unit, how much would it produce? It would produce 50 units. (round your answer to two decimal places) If the firm produced this output level, what would be the profit? its profit would be (round your answer to the nearest penny)

If the firm produced at this output level, what would the profit be? (Round to two decimal points)
Answer #1

Answer : As profit Maximisation point lies where

MR=MC

R=100q

MR=100

MC= 1.28q+25

q= 58.59 units or 59 units

For profit Maximisation level of output it would produce 59 units

Profit =TR-TC = (59*100)- ( 0.64(59)2+25(59)+1600)

Profit= 5900- ( 2227.84+1475+1600)

Profit= 5900- ( 5302.84)= $597.16